6. Two circular coil turns are so arranged in two mutually perpendicular planes that the centres of the turns coincide. The radius of each coil is 2 sm and the currents flowing through the turns I1= I2 =5 A. Find the intensity of the magnetic field at the centre of these turns.


1
Expert's answer
2022-01-30T13:52:55-0500

Answer

Current in coil I1=I2=5A

Radius of each coil R=2cm

Magnetic field due to first coil

"B'=\frac{\mu_0 I_1}{2R}"

Due to second coil

"B''=\frac{\mu_0 I_1}{2R}"

Therefore magnitude of magnetic field at centre

"B=\sqrt{B'^2+B''^2}\\=\frac{\mu_0I}{\sqrt{2}R}"

Putting all values

"B=\frac{4\times3.14\times100^{-7}\times5}{1.414\times0.02}"

B=0.02T